A sample of size 10 drawn from a normal population has a mean 31 and a variance 2.25. Is it reasonable to assume that the mean of the population is 30? 

(Use 1% level of significance, given that P(|t| )=0.1, for 9 d.f)


1
Expert's answer
2021-01-11T18:54:05-0500

Null hypothesis "H_0:\mu=30."

Alternative hypothesis "H_a:\mu\ne30."

Test statistic: "t=\frac{\bar x-mu}{\frac{s}{\sqrt{n}}}=\frac{31-30}{\frac{\sqrt{2.25}}{\sqrt{10}}}=2.11."

P-value: "p=0.0641."

Since the P-value is greater than 0.01, fail to reject the null hypothesis.

There is no sufficient evidence that the mean of of population does not equal 30.